OpenStudy (anonymous):

What is the equation of a line with a slope of 3 that goes through the point (0, -4) in slope-intercept form?

OpenStudy (midhun.madhu1987):

The point-slope form of a line is given by: y - y1= m( x - x1) where m= Slope. (x1,y1) = (0,-4) Plug in these values and simplify it into the form y = mx + c which is the slope intercept form...